IP A charge of 22.0 μC  is held fixed at the origin.

A) If a -7.00 μC charge with a mass of 3.50 g is released from rest at the position (0.925 m, 1.17 m), what is its speed when it is halfway to the origin?   

B) Suppose the -7.00 μC charge is released from rest at the point x = 12(0.925m) and y = 12(1.17m). When it is halfway to the origin, is its speed greater than, less than, or equal to the speed found in part A?

C) Find the speed of the charge for the situation described in part B.   
Answer #1

Distance between the charges,

r = sqrt ( 0.925^2 + 1.17^2 )

= 1.49 m

Potential energy of the second charge w.r.t. the first charge, kqq' / r gets fully converted into kinetic energy, (1/2) mv^2 at an infinite distance.

kqq'/r = (1/2) mv^2

v = sqrt (2kqq' / ( r*m )

= sqrt ( (2*9*10^9*7*10^(-6)*22*10^(-6))/ (1.49*0.0035) )

= 23.05 m/s